OpenStudy (anonymous):

Find the product of (x − 3)^2. x2 + 6x + 9 x2 − 9 x2 − 6x + 9 x2 + 9

OpenStudy (anonymous):

(x-3)^=(x-3)(x-3)

OpenStudy (anonymous):

try for (x-3)(x-3)

OpenStudy (asnaseer):

this site might be of help to you: http://www.algebrahelp.com/lessons/simplifying/foilmethod/

OpenStudy (anonymous):

what do i do?

OpenStudy (anonymous):

okay @asnaseer

Parth (parthkohli):

Don't use FOIL, I've got a method which has the same concept. \(\Rightarrow (x - 3)(x - 3) = x(x - 3) - 3(x - 3) \) Might help if you are familiar with distributive property.

OpenStudy (asnaseer):

@ParthKohli that method is also described on the link I gave - page 3

Parth (parthkohli):

Too lazy to open links ;)

OpenStudy (diyadiya):

You can even memorize this identity \[\Large (a-b)^2 = (a)^2 -2(a)(b) + (b)^2\]

OpenStudy (anonymous):

lol @ParthKohli I did open it..im on it now :)

OpenStudy (anonymous):

@Luis_Rivera I wrote you a testimonial after I fanned you.

OpenStudy (anonymous):

x^2-6x+9 is correct... by using (a-b)^2=a^2-2ab+b^2 where a=1,b=6 and c=9.

OpenStudy (anonymous):

@zakiaboo what is medal? Are these good for us?

OpenStudy (anonymous):

lol yes..yes it is :) It means I gave you an award for helping me! and the award is the medal.